OpenStudy (anonymous):

Use the graph below to answer this question: graph of line going through (-1,4) and (2,0)

OpenStudy (anonymous):

Find the average rate of change for the given function from x = −1 to x = 2. four-thirds negative four-thirds negative three-fourths three-fourths

OpenStudy (igreen):

Formula for Average Rate of Change: \(m = \Large\frac{y2-y1}{x2-x1}\)

OpenStudy (anonymous):

What

OpenStudy (igreen):

\(m = \Large\frac{y2-y1}{x2-x1}\) Points: \((-1, 4)\), \((2,0)\). \(m = \Large\frac{0-4}{2 + 1}\) \(m = \Large\frac{-4}{3}\)

OpenStudy (anonymous):

thankyou